# Optimising the Processing and Storage of Visibilities using lossy compression

### v1

@@ 3.5 Effect of compression on spectral line cubes
To test the impact of the compression on the well-known ‘bleeding through’ of the strong spectral feature into the surrounding channels we added this simple sky point source to one channel of the visibilities then compressed the whole dataset and imaged the data, channel by channel. Figure 10 plots the peak flux in the surrounding channels, normalised by the peak flux of the spectral feature (28.5Jy/beam). Even for a relative error bound below 10-1 the fractional error is the order of a percent, for error bounds of $10^{-2}$ and below the fractional error is about $10^{-4}$. We note that our implementation in ASKAPSoft, as it performs the MGARD compression on a (independent) channel by channel basis, will not suffer from this effect.
@@ 4.1 Conclusions
Selecting relative error bounds less than 10%, with a compression ratio of 20, introduced no significant errors in the continuum imaging, whereas 1% error bound, with a compression ratio of about 8, introduced a loss of precision about an order of magnitude less than the noise. Below 0.1% (compression ratio of 4) imperfections in the deconvolution dominated. For the spectral line imaging, EBs below 10% limited the bleed through to 2% or less of the peak flux, whereas a 1% EB limited the bleed through to less than 0.03% in the worst case. In summary, we find that EBs between $10^{-3}$ and $10^{-2}$ have a good compromise between compression and impact on the science data products.
@@ 4.2 Outlook
MGARD offers a host of additional features which may have great applications in Radio Astronomy. It continues to be improved in performance and, as it is now being optimised specifically for Radio Astronomy data, we expect further increases in compression ratios achieved. As MGARD is embedded in ADIOS2 one also has direct access to highly parallelise reading and writing, reducing I/O requirements.
@@
This research has used the facilities of the Pawsey Supercomputing Research Centre, The Oak Ridge National laboritory, the LOFAR archive and the OSKAR software. Pawsey would like to acknowledge the Whadjuk people of the Noongar nation as the traditional custodians of this country, where the Pawsey Supercomputing Research Centre is located. Pawsey pays our respects to Noongar elders past, present, and emerging. PJE and AW would like to acknowledge the many invaluable conversations with the ASKAPSOFT development team.
@@
Our Data Quality Analysis alorithms are publically avaliable on github.com/ICRAR/Image-DQA/. The LOFAR data was taken from the long-term archive lta.lofar.eu/.

### v2

@@ 3.5 Effect of compression on spectral line cubes
To test the impact of the compression on the well-known ‘bleeding through’ of the strong spectral feature into the surrounding channels we added this simple sky point source to one channel of the visibilities then compressed the whole dataset and imaged the data, channel by channel. Figure 10 plots the peak flux in the surrounding channels, normalised by the peak flux of the spectral feature (28.5Jy/beam). Even for a relative error bound below 10-1 the fractional error is the order of a percent, for error bounds of $10^{-2}$ and below the fractional error is about $10^{-4}$. We note that our implementation in ASKAPSoft, as it performs the MGARD compression in parallel on a (independent) channel by channel basis, will not suffer from this effect. However, these results will impact other future potential implementations of MGARD and thus are included for completeness.
@@ 4.1 Conclusions
Selecting relative error bounds less than 10%, with a compression ratio of 20, introduced no significant errors in the continuum imaging, whereas 1% error bound, with a compression ratio of about 8, introduced a loss of precision about an order of magnitude less than the noise. Below 0.1% (compression ratio of 4) imperfections in the deconvolution dominated. For the spectral line imaging, EBs below 10% limited the bleed through to 2% or less of the peak flux, whereas a 1% EB limited the bleed through to less than 0.03% in the worst case. In summary, we find that EBs between $10^{-3}$ and $10^{-2}$ have a good compromise between compression and impact on the science data products. High compression ratios achieved with small information loss suggest that the raw data was significantly oversampled. This is consistent with the results we have obtained, as in the signal-dominated case a smoothly changing signal is highly compressible using the recommended error bounds (i.e. 1% of the range). In the noise-dominated case we obtain more moderate compression but still up to an order of magnitude of improvement over the non-compressed data. This is because there was no useful information significantly below the resolution imposed by the thermal noise.
@@ 4.2 Outlook
MGARD offers a host of additional features which may have great applications in Radio Astronomy. It continues to be improved in performance and, as it is now being optimised specifically for Radio Astronomy data, we expect further increases in compression ratios achieved. The innovative non-linear quantisation as used by DYSCO can be implemented in MGARD to reduced the image reconstruction errors for a given compression level. As MGARD is embedded in ADIOS2 one also has direct access to highly parallelise reading and writing, reducing I/O requirements. Recent studies (Gong:2023) demonstrate that MGARD achieves a throughput of 15GB/s and 30GB/s when compressing data using an NVIDIA A100 and AMD MI250X GPU. An evaluation by Chen:2025, using 1,024 nodes on Frontier supercomputer show that MGARD can accelerate the write and read operation of ADIOS2 by 6.815.3×6.8-15.3$\times$6.8 - 15.3 × and 5.29.3×5.2-9.3$\times$5.2 - 9.3 ×, with 142379×14-2379$\times$14 - 2379 × compression ratios. Our next paper (williamson-25) discuss the performance on MGARD in ASKAPSoft on DINGO in detail.
@@
AW acknowledge the Pawsey Centre for Extreme Scale Readiness (PaCER) for funding and support. This research has used the facilities of the Pawsey Supercomputing Research Centre, The Oak Ridge National laboratory, the LOFAR archive and the OSKAR software. We would like to acknowledge the Whadjuk people of the Noongar nation as the traditional custodians of this country, where the Pawsey Supercomputing Research Centre and ICRAR is located. We pay our respects to Noongar elders past, present, and emerging. PJE and AW would like to acknowledge the many invaluable conversations with the ASKAPsoft development team.
@@
Our Data Quality Analysis algorithms are publicly available on github.com/ICRAR/Image-DQA/. The LOFAR data was taken from the long-term archive lta.lofar.eu/.